OpenStudy (anonymous):

Integrate ( (x+5) / (5-x)) ^1/2 dx How???

OpenStudy (isaiah.feynman):

The problem was explained there!

OpenStudy (anonymous):

@Isaiah.Feynman, not the same problem \[\int\sqrt{\frac{x+5}{5-x}}~dx\] Let \(u=5-x\), then \(-du=dx\). \[-\int\sqrt{\frac{10-u}{u}}~du\\ -\int\sqrt{\frac{10}{u}-1}~du\] Let \(t=\dfrac{10}{u}\), then \(-dt=\dfrac{10}{u^2}~du\), or \(-\dfrac{10}{t^2}dt=du\). \[10\int\frac{\sqrt{t-1}}{t^2}~dt\] Integrate by parts, setting \[\begin{matrix} a=\sqrt{t-1}&&&db=\frac{dt}{t^2}\\ da=\frac{dt}{2\sqrt{t-1}}&&&b=-\frac{1}{t} \end{matrix}\] This gives you \[10\left(-\frac{\sqrt{t-1}}{t}+\frac{1}{2}\int\frac{dt}{t\sqrt{t-1}}\right)\] For the integral, let \(s=\sqrt{t-1}\), which gives \(s^2=t-1\) and \(2s~ds=dt\). \[10\left(-\frac{\sqrt{t-1}}{t}+\frac{1}{2}\int\frac{2s}{(s^2+1)s}~ds\right)\\ 10\left(-\frac{\sqrt{t-1}}{t}+\int\frac{ds}{s^2+1}\right)\]